OpenStudy (anonymous):

Evaluate: 2Log_{3}12 - 2Log_{3}4 I'm going wrong somewhere, please help!

OpenStudy (anonymous):

\[2Log_{3}12 - 2Log_{3}4\] easier to read xD

OpenStudy (anonymous):

2log_{3} 12 - 2 log_{3} 4 = 2 log_{3} (12/4) = 2 log_{3} 3 = 2(1) = 2

OpenStudy (anonymous):

\[\log a - \log b = \log \frac{a}{b}\]